OpenStudy (anonymous):

can you help me out please ASAP For several years, Graham has gathered data on students who own a computer and on their performance on a typing-speed test. He creates the linear model y = 3.8x + 17.4 where y represents the typing speed, in words per minute, of a student and x represents the number of years the student has owned a computer. According to Graham’s model, what is the approximate speed of a student who has owned a computer for three years? 21 words/min 29 words/min 56 words/min 64 words/min

OpenStudy (anonymous):

I'll do the first, or try that is... y=3.8x + 17.4 21=3.8(3)+17.4 21=11.4+17.4 21=28.8 =28.8-21 =7.8 I'm not sure about it though

OpenStudy (anonymous):

ohhh, wait i think how i know how to you go, y=3.8x + 17.4 y=3.8(3)+17.4 y=11.4+17.4 y=28.8 and that's the answer, i didn't know the numbers below was options :P

OpenStudy (anonymous):

thank you very much

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

So it's 29 words/min

OpenStudy (anonymous):

thank you ;D

OpenStudy (anonymous):

no prob ;D